OpenStudy (anonymous):

Help again :p 1/5 Div 3/4=

OpenStudy (anonymous):

\[1/5\div3/4\]

OpenStudy (anonymous):

I tried the Calculator....Didn't go so well

OpenStudy (anonymous):

Lol

jhonyy9 (jhonyy9):

1/5/3/4 =1/5 *4/3 =4/15

OpenStudy (anonymous):

multiply by the reciprocal, 1/5x3/4

OpenStudy (anonymous):

1/5x4/3**

OpenStudy (ragingsquirrel):

try http://www.webmath.com/divfract.html your answer is 4/15

OpenStudy (anonymous):

0.26666666666666666666666666667

OpenStudy (anonymous):

Okay thanks :D